OpenStudy (raden):

Find all (x,y=integer number) so that x^3 - 3^y = 100

OpenStudy (anonymous):

seems unlikely do you have a solution?

OpenStudy (anonymous):

Possible typo?

OpenStudy (anonymous):

i hope so

OpenStudy (raden):

one of solution i got x=7, y=7

OpenStudy (anonymous):

otherwise you get \(y=\frac{\ln(x^3-100)}{\ln(3)}\) and

OpenStudy (raden):

sorry, x=7, y=5

OpenStudy (anonymous):

\(7^3-3^7=-1844

OpenStudy (anonymous):

i'll be damned

OpenStudy (experimentx):

x^3 = 100 + 3^y <-- exponential ^ cubic

OpenStudy (anonymous):

I'd probably go brute force on this one. *opens spreadsheet software*

OpenStudy (raden):

experimentX, yeps... exponential any one help me find all solutions, and how? i got the answer with trial and error, very tired for me...

OpenStudy (experimentx):

hold on .. they are different variables.

OpenStudy (raden):

yea, it's the problem...

OpenStudy (anonymous):

You can use the formula satellite derived and a graphing calculator or computer spreadsheet and do trial and error for numerous integer values of x and see which y values it returns are integers.

OpenStudy (anonymous):

The next one I found was (81,12) and I have a couple more after that.

OpenStudy (anonymous):

There appears to be a pattern. My conjecture is that besides the (7,5), the y values are multiples of 3.

OpenStudy (anonymous):

Furthermore, I would conjecture that the x's are integer powers of 3.

OpenStudy (anonymous):

is there one with \(x=27\) ?

OpenStudy (anonymous):

Hmmm, I seem to be getting round-off error, though . . .

OpenStudy (anonymous):

No, so far I have (7,5), (81,12), (243,15), (729, 18), (2187, 21), but after that things are looking fishy.

OpenStudy (anonymous):

Going to try to up the precision...

OpenStudy (experimentx):

(81,12) = 0

OpenStudy (anonymous):

i am trying to think what conditions make \(x^3+100\) a power of 3

OpenStudy (anonymous):

rather \(x^3-100\)

OpenStudy (anonymous):

Yep, I'm getting serious round-off error here. Drat.

OpenStudy (anonymous):

is it factorable : x^3+100 ?

OpenStudy (anonymous):

243, 15 doesn't work

OpenStudy (raden):

i think can't be factorise it mukushla...

OpenStudy (anonymous):

@CliffSedge you are getting 0 for each, not 100

OpenStudy (experimentx):

they are all zero ...

OpenStudy (anonymous):

Yep, I'm using more powerful tool now. So far I've tested all values up to x=2600 and haven't found any others.

OpenStudy (raden):

so, just one solution like i got?

OpenStudy (anonymous):

Good possibility. I'm up to x=6000 now and still haven't found any others.

OpenStudy (raden):

Ok, thanks all for ur respons dont forget, give m a medals, hahaha....

OpenStudy (anonymous):

Interesting little equation you found there. I'm going to stop looking. I'm up to x=10,000 now and still no others.

OpenStudy (experimentx):

this is a one hell of a question ... i googled.

OpenStudy (experimentx):

probably not my stufff there is a link that might lead you to solution http://math.stackexchange.com/questions/43326/on-natural-solutions-of-the-equation-y3-3x-100

OpenStudy (raden):

Ok, thank you verymuch experimentX...

OpenStudy (experimentx):

no probs at all

OpenStudy (anonymous):

tough one

OpenStudy (experimentx):

up to you man!!

OpenStudy (experimentx):

maybe someday i'll try to get number theory

OpenStudy (anonymous):

I think I found some other ones. I'm double-checking them now. I might be beyond the limit of precision of my computer. *eek*

OpenStudy (anonymous):

Dang it, I think it's still just round-off error again... Grr. Need a better processor than what I have to try brute force methods. I thought I found one around x=531441, but I'm getting x^3 - 3^y = 0 again.